OpenStudy (anonymous):

How many different batting orders are possible if 9 people are chosen from a group of 12 on a baseball team? A. 108 B. 220 C. 79,833,600 D. 387,420,489 I got C, (12P7) but it seems kinda large..

OpenStudy (anonymous):

it is large for sure

OpenStudy (anonymous):

\[12\times 11\times 10\times 9\times 8\times 7\times 6\times 5\times 4\]

OpenStudy (kayley01):

I think you are right, I think C is the answer!